6. Prepare a set of level notes for the data listed. Perform a check and adjust the misclosure. Elevation of BM 7 is 852.045 make sure to answer order of levelling question please
Answer #1

Solution Assume all readings is metres (m) Guven - Elevation of B.M7 = 852-045 m. Total loop length = 1500m. Let us make a taMisclosure m= c (kjois -1500 m - 15x10 mm C = 0.012 m m = 0.012 = 0.000 309 m (1500)005 0.000309 + 154091 m=0.309 mm) Joore h